# Z-Dip: a validated generalization of the Dip Test

### v1

@@ 3.2 Simulation of Null Distributions
To estimate $\mu_{N}$ and $\sigma_{N}$, we simulate large ensembles of null samples from the uniform distribution for a grid of sample sizes $N$ ranging from small ($N=4$) to very large (N=72, [^0] N=72{,}000). The uniform distribution is used in line with Hartigan and Hartigan’s original formulation, where it was used to calibrate empirical percentage points of the dip statistic. For each simulated dataset, we compute the Dip statistic, yielding an empirical null distribution. The mean and standard deviation of this distribution provide estimates of $\mu_{N}$ and $\sigma_{N}$.
To enable efficient application, these values are stored in precomputed lookup tables. For sample sizes not explicitly simulated, we interpolate $\mu_{N}$ and $\sigma_{N}$ linearly between adjacent grid points. This precomputation and interpolation ensure that Z-Dip scores can be rapidly obtained for any sample size without repeated simulation, and the binary file containing the complete table is provided with the software accompanying this paper.
@@ 4.1 Comparison with the Classical Dip Statistic
We first compare the behavior of the raw Dip and the standardized Z-Dip across sample sizes. The classical Dip values show strong sample-size dependence: as $N$ increases, the expected Dip under the null shrinks toward zero, while its variance decreases. This makes large-sample values numerically small and difficult to interpret.
By contrast, the Z-Dip eliminates this dependence. Across all simulated sample sizes, the null distribution of the Z-Dip is centered around zero with approximately constant spread (Figure 1, panel A). This confirms that standardization removes the sample-size dependence inherent in the raw Dip statistic.
A direct comparison between Dip $p$-values and corresponding Z-Dip values (Figure 1, panel C) further demonstrates the strong monotonic relationship between the two measures, in contrast with the regular Dip statistic (Figure 1, panel B). This alignment suggests that a single Z-Dip threshold can reproduce conventional significance levels across all $N$.
Figure 1: Comparison of the classical Dip statistic and the standardized Z-Dip across sample sizes. Z-Dip distributions were obtained over 9,999 simulated uniform samples, with 1,000 iterations of Z-Dip computation for each simulated sample. (A) Density of Z-Dip scores for different sample sizes (NN), showing strong overlap of null distributions, proving that standardization removes the dependence on $N$. (B, C) Scatterplot of Dip statistic and Z-Dip values versus corresponding Dip Test $p$-values for simulated unimodal distributions, colored by sample size $N$. The strong monotonic relationship indicates that a single Z-Dip threshold consistently matches the $p<0.05$ level across all $N$.
@@ 4.2 Threshold Calibration
Under the null hypothesis of unimodality, the Z-Dip is approximately normally distributed with mean zero and unit variance. If this normal approximation were exact, a two-tailed 5% significance level would correspond to the canonical threshold $z=1.96$.
However, empirical simulations reveal that the null distribution of the Z-Dip values is slightly right-skewed (Fig. 1, panel A). Even after standardization, minor deviations from normality persist, likely due in part to the numerical estimation of $\mu_{N}$ and $\sigma_{N}$. As a result, adopting the theoretical $z=1.96$ cutoff could yield a false-positive rate marginally above the nominal 5% level.
To ensure robustness across sample sizes, we estimate the Z-Dip threshold corresponding to $p=0.05$ empirically. For each $N$, we compute Dip-based $p$-values and corresponding Z-Dip scores for a large ensemble of null samples, then identify the Z-Dip value at which $p=0.05$. Bootstrap resampling provides confidence intervals for each estimated threshold. These values are highly consistent across $N$, with upper bounds ranging narrowly around 1.97. We therefore adopt a conservative universal cutoff $z=1.975$, corresponding closely to the theoretical $1.96$ threshold while ensuring that the nominal Type I error rate is not exceeded for any sample size.
@@ 4.3 Computational Considerations
The computation of the Dip statistic, as originally defined by Hartigan and Hartigan, involves finding the maximum deviation between the empirical distribution function and the closest unimodal distribution. The algorithm proceeds by sorting the observed data, constructing candidate unimodal envelopes, and computing the maximal deviation. The sorting step dominates the computational cost, leading to an overall time complexity of $O(n\log n)$. After sorting, the iterative identification of modal intervals and the computation of cumulative deviations are linear in $n$, so the asymptotic complexity is determined by the sorting operation.
The Z-Dip retains this computational structure. Once the Dip statistic $Dip_{obs}$ is computed, the standardized Z-dip score, defined in Formula 1, is obtained by a constant number of arithmetic operations, which have $O(1)$ complexity. Therefore, the total runtime of the Z-Dip remains $O(n\log n)$.
Crucially, the Z-Dip does not require repeated Monte Carlo sampling or permutation testing to obtain significance levels. Once lookup tables for $\mu_{N}$ and $\sigma_{N}$ are precomputed, evaluating the Z-Dip for new samples becomes effectively instantaneous. The absence of computational overhead makes the Z-Dip suitable for high-throughput or large-scale applications, such as distributional analysis across thousands of subpopulations or time points.
@@ 4.4 Synthetic Testing
denote respectively the vectors of means, standard deviations, and mixing proportions of the $K$ Gaussian components, with $\pi_{k}\geq 0$ and $\sum_{k=1}^{K}\pi_{k}=1$.
The generator function allows full control over the number, separation, and spread of modes, enabling systematic testing from clearly unimodal to strongly bimodal or multimodal cases. For each configuration, we simulated multiple sample sizes and computed both the classical Dip Test $p$-values and the corresponding Z-Dip scores.
Across all simulations, the Z-Dip preserved the inferential behavior of the Dip Test, with perfect correspondence between significance decisions.
Table 1 summarizes the results across unimodal, weakly bimodal, strongly bimodal, and trimodal conditions. The specific data-generating processes were defined as follows:
  • Unimodal: ${\mu}=(-0.3,-0.3)$, ${\sigma}=(0.1,0.1)$, ${\pi}=(1,0)$
  • Weakly bimodal: ${\mu}=(-0.6,0.6)$, ${\sigma}=(0.15,0.15)$, ${\pi}=(0.7,0.3)$
  • Strongly bimodal: ${\mu}=(-0.6,0.6)$, ${\sigma}=(0.1,0.1)$, ${\pi}=(0.5,0.5)$
@@
Multimodal mixtures yielded large positive Z-Dip values and were consistently detected as such, while unimodal distributions remained centered near zero or negative values. It can also be noted that, for genuinely multimodal distributions, the Z-Dip tends to grow systematically with sample size $N$. This behavior arises because well-separated modes produce a persistent “dip” in the empirical cumulative distribution function: as $N$ increases, the separation between peaks is resolved more clearly, leading to larger Dip statistics and correspondingly higher Z-Dip scores. This behavior is further illustrated in Fig. S1 of the Supplementary Information. In contrast, unimodal distributions exhibit flat or slowly varying Z-Dip values with $N$, since noise dominates and no persistent secondary mode exists.
Table 1: Synthetic validation of the Z-Dip statistic under varying modality and sample size (1000 replications per condition).
Configuration NN Mean Z-Dip $p<0.05$ (Dip Test) $Z>1.975$ Agreement
Unimodal 50 -0.86 0.01 0 0.998
100 -1.04 0.00 0 1.000
200 -1.22 0.00 0 1.000
400 -1.41 0.00 0 1.000
500 -1.47 0.00 0 1.000
Weak bimodal 50 7.16 1.00 1 1.000
100 11.38 1.00 1 1.000
200 17.20 1.00 1 1.000
300 21.38 1.00 1 1.000
400 25.93 1.00 1 1.000
500 28.93 1.00 1 1.000
Strong bimodal 50 9.64 1.00 1 1.000
100 15.16 1.00 1 1.000
200 22.55 1.00 1 1.000
300 28.06 1.00 1 1.000
400 33.78 1.00 1 1.000
500 37.86 1.00 1 1.000
Trimodal 50 4.48 1.00 1 0.999
100 7.10 1.00 1 1.000
200 10.70 1.00 1 1.000
300 13.34 1.00 1 1.000
400 16.19 1.00 1 1.000
500 18.12 1.00 1 1.000
These synthetic results provide a controlled validation, showing that the Z-Dip behaves as theoretically expected under both null (unimodal) and alternative (multimodal) conditions.
@@ 4.5 Empirical Validation
To evaluate the empirical validity of the proposed Z-Dip statistic, we applied it to a large-scale dataset of 117,457117{,}457 observed opinion distributions. Each distribution represents the political leaning of YouTube users engaging in political video discussions. User leanings are continuous in the range [1,+1][-1,+1], where 1-1 indicates strongly left-leaning and +1+1 strongly right-leaning users. These distributions were previously constructed and described in [4], and provide a rich and heterogeneous collection of empirical distributions, enabling large-scale testing of modality detection methods. For each distribution, we computed both the classical Dip statistic and the standardized Z-Dip, using the universal threshold $z=1.975$ for significance. Despite the empirical sample being heavily skewed toward unimodal (i.e., non-multimodal) cases, the two methods produced nearly identical classifications, with an overall agreement of $99.63\%$, as shown in Fig. 2. This close correspondence confirms that the Z-Dip effectively reproduces the inferential behavior of the original Dip Test while providing a standardized, comparable measure of deviation from unimodality.
The few discrepancies (450 cases in total) occur almost exclusively at the lowest sample sizes. The median number of observations among the disagreeing cases is $N=11$, while the third quartile is $N=52$. This pattern seems consistent with small-sample limitations of the original Dip Test: when $N$ is small, the empirical distribution contains too few distinct values to approximate a smooth shape, leading to discrete artifacts and unstable estimates. Under these conditions, even minimal numerical variation, such as rounding differences or the interpolation of $\mu_{N}$ and $\sigma_{N}$ in the Z-Dip, can flip the classification outcome.
It is also worth noting that a fixed universal threshold ($z=1.975$) necessarily ignores the slight $N$-dependence of the Dip’s true critical values. Minor disagreements are therefore expected in boundary cases, particularly for very small samples. Nonetheless, across realistic sample sizes, the agreement between the Z-Dip and the classical test is effectively perfect.
In sum, these results demonstrate that the Z-Dip not only preserves the decision structure of the Hartigan Dip Test but also generalizes it into a form that is computationally efficient, interpretable, and robust across datasets of varying sizes. In practice, the Z-Dip can be treated as a drop-in replacement for the classical Dip Test that “just works”, eliminating the need for sample-size-specific calibration without sacrificing accuracy or interpretability.
Figure 2: Confusion matrix of the agreement between the Z-Dip threshold and the Dip Test’s $p$-value over the 117,457117{,}457 empirical samples considered.
@@ 5 Addressing Large-Sample Sensitivity
A limitation of the classical Dip Test, and by extension the Z-Dip, is its increasing sensitivity at very large sample sizes. As $N$ grows, the null distribution of the Dip statistic contracts toward zero, so even minor irregularities (such as slight skewness or minimal outliers) can produce statistically significant deviations. This leads to a proliferation of “false positives”, where visually unimodal distributions are classified as multimodal. While the Z-Dip standardization improves comparability across sample sizes, the fundamental sensitivity to tiny deviations in large-$N$ samples remains.
One straightforward solution is to resample or downsample the data to a moderate effective sample size before computing the Dip statistic or the corresponding Z-Dip. By reducing $N$ to a size where the null distribution retains meaningful variance, the test becomes less sensitive to trivial deviations while preserving the ability to detect substantive multimodality. Concretely, for a large sample of size $N$, one can compute the Z-Dip for $n_{sim}$ times, with $n_{sim}$ large, on random subsets of size $N_{\text{sub}}\ll N$, and then average the resulting Z-Dip scores:
@@
where $x_{i}$ is a random subset of size $N_{\text{sub}}$, and $Z\text{-Dip}(x_{i})$ is the Z-Dip test computed on the subset $x_{i}$. This procedure preserves the nonparametric, tuning-free nature of the test, while mitigating the pathological sensitivity of the statistic in very large samples.
While marginally more computationally intensive, this downsampling-based approach enables meaningful comparisons across very large datasets and ensures that the test captures substantive deviations from unimodality rather than trivial sampling noise.
Figure 3 illustrates this effect: for two visually similar distributions, one with N=1, [^0] N=1{,}000 and another with N=100, [^0] N=100{,}000, with both having a negligible secondary peak further from the mode, the classical Dip Test, along with its $p$-value and the corresponding Z-Dip score, classify the smaller sample as unimodal, and the larger sample as multimodal. In contrast, the downsampled Z-Dip correctly identifies both distributions as unimodal ($z<1.975)$, demonstrating that it effectively controls for scale-dependent artifacts and preserves meaningful inference about the true modality of the data.
@@
Figure 4: Stable trend of downsampled Z-Dip for sample sizes $N$ ranging from 150 to 72,000, computed for data drawn from three distributions: (i) a bimodal Gaussian mixture with $\mu=(-0.6,0.6)$, $\sigma=(0.1,0.1)$, and equal component weights; (ii) a weakly bimodal distribution with $\mu=(-0.6,0.6)$, $\sigma=(0.1,0.2)$, and a minor mode weight of $0.025$; and (iii) a uniform distribution on $[0,1]$. Each point represents the mean downsampled Z-Dip over $n_{sim}=30$.
In principle, other strategies could also be applied, such as smoothing the ECDF before computing the Dip, or incorporating a sample-size-adjusted penalty term. We leave a systematic exploration of these extensions for future work, but note that even the simple downsampling approach is sufficient to restore interpretability in ambiguous, high-$N$ scenarios.

### v2

@@ 3.3 Simulation of Null Distributions
To estimate $\mu_{N}$ and $\sigma_{N}$, we simulate large ensembles ($20,000$ iterations each) of null samples from the uniform distribution for a grid of 183 roughly log-spaced sample sizes $N$, ranging from small ($N=4$) to very large (N=72, [^0] N=72{,}000). The uniform distribution is used in line with Hartigan and Hartigan’s original formulation, where it was used to calibrate empirical percentage points of the Dip statistic. For each simulated dataset, we compute the Dip statistic, yielding an empirical null distribution, and the mean and standard deviation of this distribution provide estimates of $\mu_{N}$ and $\sigma_{N}$.
For sample sizes not explicitly simulated, $\mu_{N}$ and $\sigma_{N}$ are obtained by linear interpolation between adjacent grid points, and, to enable efficient application, these values are stored in precomputed lookup tables, ensuring that Z-Dip scores can be rapidly obtained for any sample size without repeated simulation. The binary files containing the complete lookup tables are provided with the software accompanying this paper (see Section Data and code availability).
@@ 4.1 Comparison with the Classical Dip Statistic
We first compare the behavior of the raw Dip and the standardized Z-Dip across sample sizes. The classical Dip values show strong sample-size dependence: as $N$ increases, the expected Dip under the null shrinks toward zero, while its variance decreases, making large-sample values numerically small and difficult to interpret. By contrast, the Z-Dip eliminates this dependence. Across all simulated sample sizes, the null distribution of the Z-Dip is centered around zero with approximately constant spread (Figure 2, panel A). This confirms that standardization removes the sample-size dependence inherent in the raw Dip statistic.
A direct comparison between Dip $p$-values and corresponding Z-Dip scores (Figure 2, panel B) further demonstrates the strong monotonic relationship between the two measures. This alignment suggests that a single Z-Dip threshold can reproduce conventional significance levels across all $N$, in contrast with the regular Dip statistic (Figure 2, panel C).
Figure 2: Comparison of the classical Dip statistic and the standardized Z-Dip across sample sizes. Z-Dip distributions were obtained over 20,000 simulated uniform samples. (A) Density of Z-Dip scores for different sample sizes (NN), showing strong overlap of null distributions, proving that standardization removes the dependence on $N$. (B, C) Lineplot of Z-Dip and Dip statistic values versus corresponding Dip Test $p$-values for simulated unimodal distributions, colored by sample size $N$.
@@ 4.2 Threshold Calibration
Under the null hypothesis of unimodality, the Z-Dip is approximately normally distributed with mean zero and unit variance. If this normal approximation were exact, a two-tailed 5% significance level would correspond to the canonical threshold $z=1.96$. However, empirical simulations reveal that the null distribution of the Z-Dip values is slightly left-skewed (Fig. 2, panel A). Even after standardization, which enforces centering and variance normalization, but does not impose Gaussianity, minor deviations from normality persist, likely due to the inherent skewed nature of the classical Dip statistic. As a result, adopting the theoretical $z=1.96$ cutoff could yield false-negative rates marginally above the 5% level.
To ensure robustness across sample sizes, we estimate the Z-Dip threshold corresponding to $p=0.05$ empirically. For each $N\geq 20$, we compute Dip $p$-values and corresponding Z-Dip scores for a large ensemble of null samples, then identify the Z-Dip value at which $p=0.05$ via linear interpolation. Bootstrap resampling provides confidence intervals for each estimated threshold. These values are highly consistent across $N$, with upper bounds ranging narrowly around $1.85$, indicating that the residual dependence of the standardized statistic on sample size is minor. We therefore adopt a conservative universal cutoff $z=1.85$, corresponding closely to the theoretical $1.96$ threshold while ensuring that the nominal Type I error rate is not exceeded for any sample size. The choice to calibrate the threshold only on samples having $N\geq 20$ is due to the greater variability of the scores at lower sample sizes, where the coarser approximation of the empirical distribution and the reduced resolution of the Dip statistic lead to more unstable Z-Dip values and less reliable threshold calibration.
@@ 4.3 Computational Considerations
The computation of the Dip statistic, as originally defined by Hartigan and Hartigan, involves finding the maximum deviation between the empirical distribution function and the closest unimodal distribution. The algorithm proceeds by sorting the observed data, constructing candidate unimodal envelopes, and computing the maximal deviation. The sorting step dominates the computational cost, leading to an overall time complexity of $O(n\log n)$, where $n$ is the number of items to be sorted. After sorting, the iterative identification of modal intervals and the computation of cumulative deviations are linear in $n$, so the asymptotic complexity is determined by the sorting operation.
The Z-Dip retains this computational structure. Once the Dip statistic $Dip_{obs}$ is computed, the standardized Z-dip score, defined in Formula (1), is obtained by a constant number of arithmetic operations, which have $O(1)$ complexity. Therefore, the total runtime of the Z-Dip remains $O(n\log n)$. Crucially, the Z-Dip does not require repeated Monte Carlo sampling or permutation testing to obtain significance levels. Once lookup tables for $\mu_{N}$ and $\sigma_{N}$ are precomputed, evaluating the Z-Dip for new samples becomes effectively instantaneous, making it suitable for large-scale applications, such as distributional analysis across thousands of distributions [21].
@@ 4.4 Synthetic Testing
denote respectively the vectors of means, standard deviations, and mixing proportions of the $K$ Gaussian components, with $\pi_{k}\geq 0$ $\forall\,k$ and $\sum_{k=1}^{K}\pi_{k}=1$. The generator function allows full control over the number, separation, and spread of modes, enabling testing from clearly unimodal to strongly bimodal or multimodal cases. For each configuration, we simulated 280 sample sizes $N$, logarithmically spaced between 20 and 5,000, and for each $N$ performed 1,000 independent replications. For each replication, we computed both the classical Dip Test $p$-values and the corresponding Z-Dip scores.
Across all simulations, the Z-Dip preserved the inferential behavior of the Dip Test, with a nearly perfect correspondence between significance decisions: across 1,120,000 samples, only in 110 cases the two measures were not aligned in their classification, a discrepancy easily attributable to numerical approximations and/or boundary cases. The specific data-generating processes were defined as follows:
  • Unimodal: ${\mu}=(0)$, ${\sigma}=(0.1)$
  • Weak bimodal: ${\mu}=(-0.6,0.6)$, ${\sigma}=(0.15,0.15)$, ${\pi}=(0.7,0.3)$
  • Strong bimodal: ${\mu}=(-0.6,0.6)$, ${\sigma}=(0.1,0.1)$, ${\pi}=(0.5,0.5)$
@@
Note how the unimodal case simply represents samples drawn from a normal distribution, and that in the in weak bimodal case, we provide $\sigma_{1}=\sigma_{2}=0.15$, instead of $\sigma=0.1$ used in the other cases, to make the modes slightly less peaked and prominent.
Table 1 summarizes the results across unimodal, weak bimodal, strong bimodal, and trimodal conditions. Multimodal mixtures yielded large positive Z-Dip values and were consistently detected as such, while unimodal distributions remained centered near zero or negative values. It can also be noted that, for genuinely multimodal distributions, the Z-Dip tends to grow systematically with sample size $N$. This behavior arises because well-separated modes produce a persistent “dip” in the empirical cumulative distribution function: as $N$ increases, the separation between peaks is resolved more clearly, leading to larger Dip statistics and correspondingly higher Z-Dip scores. This behavior is further illustrated in Fig. S2 in the SI. In contrast, unimodal distributions exhibit flat or slowly varying Z-Dip values with $N$, since noise dominates and no persistent secondary mode exists. These synthetic results provide a controlled validation, showing that the Z-Dip behaves as theoretically expected under both null (unimodal) and alternative (multimodal) conditions.
Table 1: Synthetic validation of the Z-Dip statistic under varying modality and sample size (1,000 replications per each combination of $N$ and configuration).
Configuration $N$ (range) Mean Z-Dip $p<0.05$ (Dip Test) $Z>1.85$ # of simulations Agreement Unimodal 20-50 -0.750 0.003 0.003 31,000 1.000 51-100 -0.952 0.000 0.000 37,000 1.000 101-500 -1.253 0.000 0.000 87,000 1.000 501-1000 -1.547 0.000 0.000 37,000 1.000 >>1000 -1.835 0.000 0.000 88,000 1.000 Weak bimodal 20-50 5.515 0.971 0.971 31,000 0.999 51-100 9.180 1.000 1.000 37,000 1.000 101-500 19.000 1.000 1.000 87,000 1.000 501-1000 35.434 1.000 1.000 37,000 1.000 >>1000 67.489 1.000 1.000 88,000 1.000 Strong bimodal 20-50 7.517 0.994 0.994 31,000 1.000 51-100 12.320 1.000 1.000 37,000 1.000 101-500 25.001 1.000 1.000 87,000 1.000 501-1000 46.149 1.000 1.000 37,000 1.000 >>1000 87.221 1.000 1.000 88,000 1.000 Trimodal 20-50 3.459 0.900 0.897 31,000 0.996 51-100 5.724 0.999 0.999 37,000 1.000 101-500 11.850 1.000 1.000 87,000 1.000 501-1000 22.186 1.000 1.000 37,000 1.000 >>1000 42.414 1.000 1.000 88,000 1.000
@@ 4.5 Empirical Validation
To evaluate the empirical validity of the proposed Z-Dip statistic, we applied it to a large-scale dataset of 88,52888{,}528 observed opinion distributions having sample size $N\geq 20$. Each distribution represents the political leaning of YouTube users engaging in political video discussions. User leanings are continuous in the range [1,+1][-1,+1], where 1-1 indicates strongly left-leaning and +1+1 strongly right-leaning users. These distributions were previously constructed and described in [21], and provide a rich and heterogeneous collection of empirical distributions, enabling large-scale testing of modality detection methods. For each distribution, we computed both the classical Dip statistic and the standardized Z-Dip, using the threshold $z=1.85$ for significance derived in Section 3.
Despite the empirical sample being heavily skewed toward unimodal (i.e., non-multimodal) cases, the two methods produced nearly identical classifications, with an overall agreement of $99.899\%$. Moreover, the two measures yield a Cohen’s Kappa coefficient of $\kappa=0.981$, indicating near-perfect agreement that is well above what would be expected by chance. This close correspondence confirms that the Z-Dip effectively reproduces the inferential behavior of the original Dip Test while providing a standardized, comparable measure of deviation from unimodality. Fig. 3 further illustrates these results via a column-normalized confusion matrix between the two measures.
The few discrepancies (89 cases in total) often occur at lower sample sizes. The median number of observations among the disagreeing cases is $N=89$. This likely stems from the fact that when $N$ is small, the empirical distribution contains too few distinct values to approximate a smooth shape, leading to discrete artifacts and unstable estimates. Under these conditions, small numerical variation, such as rounding differences or the interpolation of $\mu_{N}$ and $\sigma_{N}$ in the Z-Dip, can flip the classification outcome. It is also worth noting that a fixed universal threshold ($z=1.85$) necessarily ignores the slight $N$-dependence of the Dip’s true critical values. Minor disagreements are therefore expected in boundary cases. Nonetheless, across our empirical validation, the agreement between the Z-Dip and the classical test is essentially perfect.
In summary, the Z-Dip preserves the decision structure of Hartigans’ Dip Test and generalizes it into a form that is computationally efficient, interpretable, and robust across datasets of varying sizes. Furthermore, the Z-Dip can be treated as a drop-in replacement for the classical Dip Test that eliminates the need for sample-size-specific calibration without sacrificing accuracy or interpretability.
Figure 3: Column-normalized confusion matrix of the agreement between the Z-Dip threshold and the Dip Test’s $p$-value over the 88,52888{,}528 empirical samples considered.
@@ 5 Addressing Large-Sample Sensitivity
A limitation of the classical Dip Test, and by extension the Z-Dip, is its increasing sensitivity at very large sample sizes. As $N$ grows, the null distribution of the Dip statistic contracts toward zero, so even minor irregularities (such as slight skewness or minimal outliers) can produce statistically significant deviations. This leads to a proliferation of practically negligible deviations that are treated as significant, where effectively unimodal distributions are classified as multimodal.
While the Z-Dip standardization improves comparability across sample sizes, the fundamental sensitivity to tiny deviations in large-$N$ samples remains.
One straightforward solution is to resample or downsample the data to a moderate effective sample size before computing the Dip statistic or the corresponding Z-Dip. By reducing $N$ to a size where the null distribution retains meaningful variance, the test becomes less sensitive to trivial deviations while preserving the ability to detect substantive multimodality. Concretely, for a large sample of size $N$, one can compute the Z-Dip for $n_{sim}$ times on random subsets of size $N_{\text{sub}}\ll N$, and then average the resulting Z-Dip scores:
@@
where $x_{i}$ is a random subset of size $N_{\text{sub}}$, and $Z\text{-Dip}(x_{i})$ is the Z-Dip score computed on the subset $x_{i}$. This procedure preserves the nonparametric, tuning-free nature of the test, while mitigating the pathological sensitivity of the statistic in very large samples. This downsampling-based approach enables meaningful comparisons across very large datasets and ensures that the test captures substantive deviations from unimodality rather than trivial sampling noise, adding only very little computational overhead.
Figure 4 illustrates this effect: for two visually similar distributions, one with N=1, [^0] N=1{,}000 and another with N=100, [^0] N=100{,}000, with both having a negligible secondary peak further from the mode, the classical Dip Test, along with its $p$-value and the corresponding Z-Dip score, classify the smaller sample as unimodal, and the larger sample as multimodal. In contrast, the downsampled Z-Dip correctly identifies both distributions as unimodal ($z<1.85)$, demonstrating that it effectively controls for scale-dependent artifacts and preserves meaningful inference about the true modality of the data.
@@
Figure 5: Stable trend of downsampled Z-Dip for sample sizes NN ranging from 150 to 72,000, computed for data drawn from three distributions: (i) a bimodal Gaussian mixture with $\mu=(-0.6,0.6)$, $\sigma=(0.1,0.1)$, and equal component weights; (ii) a weak bimodal distribution with $\mu=(-0.6,0.6)$, $\sigma=(0.1,0.2)$, and a minor mode weight of $0.025$; and (iii) a uniform distribution on $[0,1]$. Each point represents the mean downsampled Z-Dip over $n_{sim}=30$.
In principle, other strategies could be applied, such as modeling the scaling parameter of the ECDF, or incorporating a sample-size-adjusted penalty term. However, we note that even this simple downsampling approach is sufficient to restore interpretability in high-$N$ scenarios.